ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 216, 2016

Abadias, L., C. Lizama, P. J. Miana and M. P. Velasco Cesàro sums and algebra homomorphisms of bounded operators

471

Akopyan, A. and A. Balitskiy Billiards in convex bodies with acute angles

833

Balitskiy, A. See A. Akopyan Burillo, J., F. Matucci and E. Ventura The conjugacy problem

in extensions of Thompson’s group F

15

Bushaw, N., K. Gunderson and S. Kalikow Random Markov processes for countable and uncountable alphabets

181

Casini, E., E. Miglierina, Ł. Piasecki and L. Veselý Rethinking polyhedrality for Lindenstrauss spaces

355

Chang, W., C. Cheng and M. Leu A remark on the ring of algebraic integers in ℚ(\sqrt{−d})

605

Cheng, C. See W. Chang Choi, D. and S. Lim Finite-dimensional period spaces for the spaces

of cusp forms

507

Cholak, P. A., D. D. Dzhafarov and M. I. Soskova Genericity for Mathias forcing over general Turing ideals

583

Dolfi, S., R. Guralnick, C. E. Praeger and P. Spiga On the maximal number of coprime subdegrees in finite primitive permutation groups

107

Donoso, S. and W. Sun A pointwise cubic average for two commuting transformations

657

Dowdall, S., I. Kapovich and S. J. Taylor Cannon–Thurston maps for hyperbolic free group extensions

753

Downey, R., N. Greenberg and J. S. Miller Generic Muchnik reducibility and presentations of fields

371

Dzhafarov, D. D. See P. A. Cholak Friedberg, S. and L. Zhang Tokuyama-type formulas for charac-

ters of type B

617

Frigerio, R., C. Löh, C. Pagliantini and R. Sauer Integral foliated simplicial volume of aspherical manifolds

707

Garrido, A. On the congruence subgroup problem for branch groups  1

Greenberg, N. See R. Downey Gunderson, K. See N. Bushaw

Gundert, A. and U. Wagner On eigenvalues of random complexes  545

Guralnick, R. See S. Dolfi Hartman, Y. and O. Tamuz Stabilizer rigidity in irreducible group

actions

679

Hong, J. Separably closed valued fields: immediate expansions  811

Hull, M. and D. Osin Transitivity degrees of countable groups and acylindrical hyperbolicity

307

Kalikow, S. See N. Bushaw Kapovich, I. See S. Dowdall Kohlenbach, U. On the quantitative asymptotic behavior of strongly

nonexpansive mappings in Banach and geodesic spaces

215

Leu, M. See W. Chang Li, H., S. Tan and Q. Wang A certain Clifford-like algebra and

quantum vertex algebras

441

Lim, S. See D. Choi Livshyts, G., G. Paouris and P. Pivovarov On sharp bounds for

marginal densities of product measures

877

Lizama, C. See L. Abadias Löh, C. See R. Frigerio Lücke, P., L. M. Ros and P. Schlicht The Hurewicz dichotomy

for generalized Baire spaces

973

Marks, G. and R. Mazurek Rings with Linearly Ordered Right Annihilators

415

Matucci, F. See J. Burillo Mazurek, R. See G. Marks Meyerovitch, T. and A. Yadin Harmonic functions of linear

growth on solvable groups

149

Miana, P. J. See L. Abadias Miglierina, E. See E. Casini Mihara, T. On Tate’s acyclicity and uniformity of Berkovich spectra

and adic spectra

61

Miller, J. S. See R. Downey

Nikolov, N. Verbal width in anabelian groups  847

Osin, D. See M. Hull Pagliantini, C. See R. Frigerio Paouris, G. See G. Livshyts Pasten, H. and X. Vidaux Positive existential definability of multi-

plication from addition and the range of a polynomial 273 Patey, L. The weakness of being cohesive, thin or free in reverse math-

ematics

905

Perles, M. A. and M. Sigron Some variations on Tverberg’s theorem

957

Piasecki, Ł See E. Casini Pivovarov, P. See G. Livshyts Praeger, C. E. See S. Dolfi Ros, L. M. See P. Lücke Sauer, R. See R. Frigerio Schlicht, P. See P. Lücke Sigron, M. See M. A. Perles

Simon, S. Average-value Tverberg partitions via finite Fourier analysis  891

Sinapova, D. The tree property at the first and double successors of a singular

799

Soskova, M. I. See P. A. Cholak Spiga, P. See S. Dolfi

Sun, J. Module twistors for modules of nonlocal vertex algebras  247

Sun, W. See S. Donoso Tamuz, O. See Y. Hartman Tan, S. See H. Li Taylor, S. J. See S. Dowdall Velasco, M. P. See L. Abadias Ventura, E. See J. Burillo Veselý, L. See E. Casini Vidaux, X. See H. Pasten Wagner, U. See A. Gundert Wang, Q. See H. Li Yadin, A. See T. Meyerovitch Yang, L. Expanding curves in T^{1}(ℍ^{n}) under geodesic flow and equidis-

tribution in homogeneous spaces

389

Zhang, L. See S. Friedberg